A van travels220 miles on 10 gallons of gas. how many gallons the van needs to travel 550 miles. I need help in writing and solving how.
Answer by stanbon(75887)

A van travels220 miles on 10 gallons of gas. how many gallons the van needs to travel 550 miles. I need help in writing and solving how.
------------
x/550=10/220
x=550(1/22)
x=25 gallons to travel 550 miles.
Cheers,
Stan H.